OpenStudy (anonymous):

Convert to % 0.078 = 7.8%??

OpenStudy (anonymous):

yes, correct

Parth (parthkohli):

Eh?

OpenStudy (anonymous):

multiply the number with 100

OpenStudy (anonymous):

Thing of the sign % as this: When converting to: x100 When it's next to a number and you need to convert to a decimal: /100

Parth (parthkohli):

0.078 = 7.8% Yes.

Parth (parthkohli):

0.078 = 78/1000 = 7.8/100 = 7.8% I hope you get it :)

OpenStudy (anonymous):

got it thanks